# How Far Can Transformers Reason? The Globality Barrier and Inductive Scratchpad

### v2

@@ 3.2 Inductive Scratchpads
As discussed previously, scratchpads can break the global reasoning barrier with appropriate mid-steps. In this part, however, we show that fully educated scratchpads can be sensitive to the number of reasoning steps, translating into poor out-of-distribution (OOD) generalization. As a remedy, we put forward the concept of inductive scratchpad which applies to a variety of reasoning tasks as in previous sections.
@@ 3.2.1 Educated scratchpad can overfit in-distribution samples
Consider the cycle task with $40$ nodes. For the test distribution, we use the normal version of the cycle task, i.e., either two cycles of size $20$ and the nodes are not connected or a single cycle of size $40$ where the distance between the query nodes is $20$. For the train distribution, we keep the same number of nodes and edges (so the model does not need to rely on new positional embeddings for the input) but break the cycles to have uneven lengths: (1) a cycle of size $10$ and a cycle of size $30$ when the two nodes are not connected (the source query node is always in the cycle of size $10$) or (2) a cycle of size $40$ where the nodes are at distance $10$. Thus, in general, we always have $40$ nodes/edges in the graphs. However, the length of the DFS path (i.e., number of reasoning steps) is $10$ at training and $20$ at test. We trained our model on this version of the task with the DFS scratchpad. The results are shown in Figure 4(b). We observe that the model quickly achieves perfect accuracy on the training distribution, yet, it fails to generalize OOD as the model overfits the scratchpad length. In the next part, we introduce the notion of inductive scratchpad to fix this problem.
@@ 3.2.2 Inductive scratchpad: definition and experimental results
The inductive scratchpad for an inductive task is similar in format to the fully educated scratchpad but it has the following modifications: (1) tokens: two new special tokens are used: the <START> token which separates the question from the intermediate states and the <STATE> token (denoted # hereafter) to separate the states. Using these tokens, for an input question $Q$, the format of the inductive scratchpad reads <START>s[1]#s[2]#...#s[k]<EOS>. (2) generation: we want the model to promote induction and thus ‘forget’ all the previous states except the last one for the new state update. I.e., we want to generate tokens of s[i+1] as the input was Q<START>s[i]#. To implement this, one can use attention masking and reindex positions (in order to have a proper induction) or simply remove the previous states at each time; (3) training: when training the scratchpad, we want the model to learn the induction function $g$, i.e., learning how to output s[i+1]# from Q<START>s[i]#, which can be achieved with attention masking and reindexing the positions. As a result, the inductive scratchpad can be easily integrated with the common language models without changing their behavior on other tasks/data. We refer to Appendix C.2 for a detailed description of the inductive scratchpad implementation.
@@
We can use inductive scratchpads to achieve length generalization for the parity and addition tasks. For parities, we insert random spaces between the bits and design an inductive scratchpad based on the position of the bits and then compute the parity iteratively. Using this scratchpad we can train a Transformer on up to $30$ bits and get generalization on up to $50$, or $55$ bits. Results are provided in Figure 5(a). For the addition task, we consider two inductive scratchpad formats. We provide an inductive scratchpad that requires random spaces between the digits in the input and uses the position of the digits to compute the addition digit-by-digit (similar to the parity). With this scratchpad, we can generalize to numbers with 20 digits while training on numbers with up to 10 digits. In the other format, we use random tokens in the input and we compute the addition digit-by-digit by shifting the operands. The latter enables us to generalize from 4 to 26 digits at the cost of having a less natural input format. The results for different seeds are provided in Figure 5(b). See details of these scratchpads in Appendices B.4B.5.121212Our code is also available at https://github.com/aryol/inductive-scratchpad. Also see Appendix A for a comparison with recent methods.
@@ 4 Conclusion
This paper shows that for the learning objective and in contrast to expressivity results, Transformers trained from scratch have a ‘global reasoning barrier’ quantified by the globality degree: if the target requires a global composition of the tokens, Transformers will not efficiently learn. The globality measure has a simpler form and broader applicability range than prior measures as discussed in Appendix A.2, it also has tighter applicability for Transformers. The measure is currently defined for weak learning (inverse-polynomial or constant edge), and a natural next step is to consider stronger learning requirements with notions of ‘globality leap’, e.g., expanding the current work in the direction of [28] but for more general distributions. Investigating the role of curriculum learning is another natural direction and we provide preliminary results here in Appendix B.2.
@@
The inductive scratchpad can be used for a broad range of reasoning/algorithmic tasks and can easily be integrated into Transformers. The inductive scratchpad is independent of the number of reasoning steps since the model only learns the induction function. Consequently, the model thus better generalizes to inputs requiring different numbers of reasoning steps. This gives improvements of OOD/length generalization for the cycle task (Figure 4(b)), parity (Figure 5(a)), and addition (Figure 5(b)).
@@ A.1 Reasoning capabilities of Transformers
Nye et al. [32] put forward the idea of using scratchpads by showing that training Transformers on the intermediate steps in addition to the final answer can improve their performance on different tasks such as executing Python code, addition, and evaluating polynomials. Similarly, in chain-of-though reasoning (CoT) [60] models are shown step-by-step demos of problem-solving (or models generate chains of thought without any examples as in zero-shot CoT [61], among other variants). It has been further shown that using explicit explanations and reducing ambiguity can be proven useful as in the notion of algorithmic prompting [62]. Lanchantin et al. [63] introduced the concept of self-notes, showing that interleaving the intermediate reasoning steps within the question/context rather than after the question can boost the performance of scratchpad/CoT on certain tasks. Goyal et al. [64] introduced pause tokens which act as dummy tokens that provide models with more compute time and processing before the generation of the true next token. On a related note, the iterative prompting method in [65] involves querying LLMs iteratively to optimize question prompts. Further, the Select and Inference framework (SI) [66] utilizes pre-trained LLMs as general processing modules, alternating between selection and inference to generate interpretable reasoning steps leading to the final results. In this work, we used the concept of autoregressive globality to explain why scratchpads are helpful and how to design them. Further, we introduced inductive scratchpads that are suitable for a broad range of algorithmic reasoning tasks and can potentially generalize to more complex samples than what appears in the train distribution.
Parity and addition.
@@
Parity and addition tasks are two essential reasoning tasks that are challenging in the setting of length generalization (where the number of bits/digits is increased). These tasks can also be hard in the in-distribution setting if the number of bits/digits is large enough. It has been shown that certain formats of scratchpad/chain-of-though reasoning and relative positional embedding can achieve a modest length generalization on parity task (around 5/10 more bits depending on the method and the seed) [11, 53]. Lee et al. [45] use a decoder-only model similar to ours (also similar in size) showing that generating the output in the reverse format and using scratchpads are helpful. However, they are not able to get length generalization with their model size. Nevertheless, it has been shown that large enough models (with more than $10^{8}$ parameters) with a scratchpad can generalize to numbers with 9/10 digits while being trained on numbers with up to 8 digits [32]. Jelassi et al. [67] show that encoder-only models with relative positional embedding can generalize to numbers with 15 digits when they are trained on numbers with up to 5 digits.
The RASP-L work [68] considers the addition task and can get length generalization from 35/40 to 50 digits (depending on the random seed) by outputting the result in the reverse order and also using ‘index hints’ which are special tokens that come before the operands’ digits in a specific order. For instance, an example would look like $a5b4+a3b7=b1a9$. Likewise, for parity targets, they use index hints along with scratchpad and can get length generalization from 30/35 bits to 50 bits (depending on the seed).
The work of Shen et al. [17] uses a scratchpad with a recursive format for the addition task to generalize to numbers with 12/13 digits while the training data includes numbers with up to 10 digits. Our special case of inductive scratchpad based on shifting the numbers is similar to their recursive scratchpad in terms of format, however, [17] does not enforce any inductive structure (as achieved with the attention masking for the inductive scratchpad) and hence [17] gets a much more limited length generalization.
In this work, we proposed inductive scratchpads for the parity and addition tasks that only require the use of random spaces in the question formatting. For parity, our scratchpad can easily generalize to 50 or 55 (depending on the seed) bits when trained on samples with up to 30 bits. The inductive scratchpad for the addition that requires random spaces in the question can also generalize to numbers with 18/20 digits when trained on numbers with up to 10 digits. We have also provided another scratchpad based on shifting the operands at each step for the addition task that can generalize to numbers with up to 26 digits when trained on numbers with up to 4 digits at the cost of having a less natural question format.
Reasoning over graphs.
@@
On the modeling front, a group of efforts have primarily focused on baking the appropriate inductive biases into the models for reasoning and logical inference over graphs. The work [76] proposes architectural innovations to use both the context and external knowledge sources or as in [77] for learning compositional logical rules. The Neurosymbolic approach, presented in [78] involves performing semantic parsing on natural language input using neural components to transform the input into discrete structures suitable for symbolic reasoning by the logic theorem provers. Graphomer [79] extends standard Transformer architecture with a graph reasoning inductive bias. The work of [54] combines Transformers with the inductive bias of the recurrent models. Other works take the route of generating more data using external sources like a Knowledge Graph [80] or extending pre-training [81]. The work in [74] evaluates LLMs’ graph reasoning capabilities in the presence of various prompting techniques like scratchpads/CoT and their variants. This study introduces Build-a-Graph Prompting, an instruction-based prompting method, as well as Algorithmic Prompting, which includes references to relevant graph algorithms in the prompts, as two approaches for enhancing LLMs in solving graph tasks in natural language.
A.2 Learning measures and lower-bounds for GD
@@ B.1 Implications on random graphs
Figure 7: Performance of a model trained on a balanced distribution of random graphs with 24 nodes and edges where with probability $0.5$ the query nodes are not connected and with probability $0.5$ they are connected and their distance is uniformly selected from $1,2,3,4$. The validation set has the same distribution as the training set showing that the model reaches around $80\%$ accuracy on in-distribution samples. Particularly, the model has perfect accuracy on connected nodes (distance 1-4) and around $60\%$ accuracy on the nodes that are not connected. However, when we tested the model on OOD samples (where some spurious correlations are not present) the model shows a chance level performance. Note that these samples would be of low complexity if the model was actually checking whether there exists a path or not.
@@
We tried to further understand the behavior of this model. By sampling, we computed that one can get an accuracy of around $82\%$ on in-distribution samples just by outputting not-connected if the out-degree of the source query node or the in-degree of the destination query node is zero and connected otherwise. Further, we noticed that this predictor has a high correlation with the output of the model. In particular, in almost all of the cases that the model predicts not-connected, the source’s out-degree or the destination’s in-degree is zero. (The model may still misclassify some of such samples depending on the random seed.) The latter shows that the model is indeed relying on the degrees of the query nodes as a shortcut.

### v3

@@ 3.2 Inductive Scratchpads
As discussed, scratchpads can break the global reasoning barrier with appropriate mid-steps. In this part, however, we show that fully educated scratchpads can be sensitive to the number of reasoning steps, translating into poor out-of-distribution (OOD) generalization. As a remedy, we put forward the concept of inductive scratchpad which applies to various reasoning tasks as in previous sections.
@@ 3.2.1 Educated scratchpad can overfit in-distribution samples
Consider the cycle task with $24$ nodes. For the test distribution, we use the normal version of the cycle task, i.e., either two cycles of size $12$ and the nodes are not connected or a single cycle of size $24$ where the distance between the query nodes is $12$. For the train distribution, we keep the same number of nodes and edges (so the model does not need to rely on new positional embeddings for the input) but break the cycles to have uneven lengths: (1) a cycle of size $6$ and a cycle of size $18$ when the two nodes are not connected (the source query node is always in the cycle of size $6$) or (2) a cycle of size $24$ where the nodes are at distance $6$. Thus, in general, we always have $24$ nodes/edges in the graphs. However, the length of the DFS path (i.e., number of reasoning steps) is $6$ at training and $12$ at test. We trained our model on this version of the task with the DFS scratchpad. The results are shown in Figure 4(b). We observe that the model quickly achieves perfect accuracy on the training distribution, yet, it fails to generalize OOD as the model overfits the scratchpad length and number of reasoning steps. In the next part, we introduce the notion of inductive scratchpad to fix this problem.
@@ 3.2.2 Inductive scratchpad: definition and experimental results
The inductive scratchpad for an inductive task is similar in format to the fully educated scratchpad but it has the following modifications: (1) tokens: two new special tokens are used: the <START> token which separates the question from the intermediate states and the <STATE> token (denoted # hereafter) to separate the states. Using these tokens, for an input question $Q$, the format of the inductive scratchpad reads <START>s[1]#s[2]#...#s[k]<EOS>. (2) generation: we want the model to promote induction and thus ‘forget’ all the previous states except the last one for the new state update. I.e., we want to generate tokens of s[i+1] as if the input was Q<START>s[i]#. To implement this, one can use attention masking and reindex positions (in order to have a proper induction) or simply remove the previous states at each time; (3) training: when training the scratchpad, we want the model to learn the induction function $g$, i.e., learning how to output s[i+1]# from Q<START>s[i]#, which can be achieved with attention masking and reindexing the positions. As a result, the inductive scratchpad can be easily integrated with the common language models without changing their behavior on other tasks/data. We refer to Appendix C.2 for a detailed description of the inductive scratchpad implementation.
@@
We can use inductive scratchpads to achieve length generalization for the parity and addition tasks. For parities, we insert random spaces between the bits and design an inductive scratchpad based on the position of the bits and then compute the parity iteratively. The performance of this inductive scratchpad is depicted in Figure 5(a) where we can see a Transformer trained on inputs with up to $30$ bits can generalize to samples with up to $50/55$ bits depending on the seed. For the addition task, we propose two inductive scratchpads. (1) Random space method that requires random spaces between the digits in the input and uses the position of the digits to compute the addition digit-by-digit (similar to the parity). With this scratchpad, we can generalize to numbers with 18 digits while training on numbers with up to 10 digits. (2) Shift method that uses random tokens in the input and computes the addition digit-by-digit by shifting the operands. The latter enables us to generalize from 4 to 26 digits at the cost of having a less natural input formatting. The results for different seeds are provided in Figure 5(b). See details of these scratchpads in Appendices B.4B.5.121212Our code is available at https://github.com/aryol/inductive-scratchpad. A rough comparison between the performance of different methods for addition is given in Table 1. Note that the settings used in these works are not exactly the same, e.g., our methods often work with smaller models and more natural input formatting. See Appendix A.2 for a detailed comparison for both the parity and addition tasks.
Table 1: Length generalization of different methods for the addition task where our methods are shown in bold. $a\rightarrow b$ means generalizing to $b$ digits when trained on $a$ digits.
Method [34] [17] [35] [36] [37] random space method shift method
Performance $8\rightarrow 9$ $10\rightarrow 12$ $40\rightarrow 50$ $5\rightarrow 15$ $40\rightarrow 65$ $10\rightarrow 18$ $4\rightarrow 26$
@@ 4 Conclusion
This paper shows that for the learning objective and in contrast to expressivity results, Transformers trained from scratch have a ‘global reasoning barrier’ quantified by the globality degree. The globality measure has a simpler form and broader applicability range than prior measures as discussed in Appendix A.4, it also has tighter applicability for Transformers. The measure is currently defined for weak learning (inverse-polynomial or constant edge), and a natural next step is to consider stronger learning requirements with notions of ‘globality leap’, e.g., expanding the current work in the direction of [28] but for more general distributions. Investigating the role of curriculum learning is another natural direction and we provide preliminary results here in Appendix B.2.
@@
The inductive scratchpad can be used for a broad range of reasoning/algorithmic tasks and can easily be integrated into Transformers. The inductive scratchpad is independent of the number of reasoning steps since the model only learns the induction function. Consequently, the model better generalizes to inputs requiring different numbers of reasoning steps. This gives improvements of OOD/length generalization for the cycle task (Figure 4(b)), parity (Figure 5(a)), and addition (Figure 5(b)).
@@ A.1 Reasoning capabilities of Transformers
Nye et al. [32] put forward the idea of using scratchpads by showing that training Transformers on the intermediate steps in addition to the final answer can improve their performance on different tasks such as executing Python code, addition, and evaluating polynomials. Similarly, in chain-of-though reasoning (CoT) [63] models are shown step-by-step demos of problem-solving (or models generate chains of thought without any examples as in zero-shot CoT [64], among other variants). It has been further shown that using explicit explanations and reducing ambiguity can be proven useful as in the notion of algorithmic prompting [65]. Lanchantin et al. [66] introduced the concept of self-notes, showing that interleaving the intermediate reasoning steps within the question/context rather than after the question can boost the performance of scratchpad/CoT on certain tasks. Goyal et al. [67] introduced pause tokens which act as dummy tokens that provide models with more compute time and processing before the generation of the true next token. On a related note, the iterative prompting method in [68] involves querying LLMs iteratively to optimize question prompts. Further, the Select and Inference framework (SI) [69] utilizes pre-trained LLMs as general processing modules, alternating between selection and inference to generate interpretable reasoning steps leading to the final results.
The work of [70] studies the role of scratchpad from an expressivity point of view. Assuming $TC^{0}\neq NC^{1}$, they show the existence of tasks that are not expressible by constant-depth Transformers without scratchpad and expressible by constant-depth Transformers with scratchpad. It is further shown that dynamic programming (DP) algorithms can be expressed by constant-depth Transformers with scratchpad [70]. In this work, however, we used the concept of autoregressive globality to explain why scratchpads are helpful and how to design them from the learning point of view. Further, we introduced inductive scratchpads that are suitable for a broad range of algorithmic reasoning tasks and can potentially generalize to more complex samples than what appears in the train distribution. We note that inductive scratchpads are also suitable for DP algorithms thanks to their iterative structure (the variables of the DP algorithm can be updated in each state of the inductive scratchpad). We stress that our inductive scratchpad potentially enables models to generalize to OOD samples for such DP algorithms and/or to work with longer scratchpads because of the attention masking which reduces the effective context size. In concurrent research, [71] shows that a restricted class of iterative algorithms with scratchpad (including parity task) are expressible by 2-layer Transformers. Compared to our inductive scratchpad, [71] does not have any length generalization result.
A.2 Length generalization for parity and addition tasks
Parity and addition tasks are two essential reasoning tasks that are challenging in the setting of length generalization where the number of bits/digits is increased. These tasks can also be hard in the in-distribution setting if the number of bits/digits is large enough. It has been shown that certain forms of scratchpad/chain-of-though reasoning and relative positional embedding can achieve a modest length generalization on the parity task (around 5/10 more bits depending on the method and the seed) [11, 34]. The RASP-L work [35] considers the parity task and can get length generalization from 30 to 50 bits (depending on the random seed) by using ‘index hints’ which are special tokens that come before the input bits in a specific order. For instance, an input could look like $a0b0c1d1e0$. Our proposed inductive scratchpad for the parity task only requires the use of random spaces in the question formatting and can generalize to 55 bits when trained on samples with up to 30 bits. Different works use different input formatting, techniques, model sizes, and train/test datasets, nevertheless, a rough comparison between different works for the parity task is given in Table 2 (learning is defined as above $80\%$ accuracy for the majority/median of the results for different seeds, some seeds do better than others).
Table 2: Length generalization performance of different methods for the parity task
Work Performance Method and Assumptions
Kazemnejad et al. [34] From 8 bits to 12 bits Using NoPE (no positional embedding)
RASP-L work [35] From 30 bits to 50 bits Using scratchpad + ‘index hints’ (special tokens before each bit in the input), an input & output look like $a0b0c1d1e0>+c-d+$
Anil et al. [11] From 8 bits to 20 bits Using large pre-trained models (128B) + prompting + fine-tuning + scratchpad
Our method From 30 bits to 55 bits Using random spaces in the input (e.g., _01_10_0__1_) + inductive scratchpad
@@ A.2 Length generalization for parity and addition tasks
For the addition task, Lee et al. [49] use a decoder-only model similar to ours (also similar in size) showing that generating the output in the reverse format and using scratchpads are helpful. However, they are not able to get length generalization with their model size. Nevertheless, it has been shown that large enough models (with more than $10^{8}$ parameters) with a scratchpad can generalize to numbers with 9/10 digits while being trained on numbers with up to 8 digits [32]. Jelassi et al. [36] show that encoder-only models with relative positional embedding can generalize to numbers with 15 digits when they are trained on numbers with up to 5 digits. The RASP-L work [35] considers the addition task and can get length generalization from 35/40 to 50 digits (depending on the seed) by outputting the result in the reverse order and also using ‘index hints’ (special tokens coming before the operands’ digits in a specific order). For instance, an example would look like $a5b4+a3b7=b1a9$. Zhou et al. [37] further improve the latter by using FIRE relative positional embedding [72] and randomized position encoding [73], generalizing to numbers with 65 digits when trained on numbers with up to 40 digits. The work of Shen et al. [17] uses a scratchpad with a recursive format for the addition task to generalize to numbers with 12/13 digits while the training data includes numbers with up to 10 digits. Our special case of inductive scratchpad based on shifting the numbers is similar to their recursive scratchpad, however, [17] does not enforce any inductive structure (as achieved with the attention masking for the inductive scratchpad) and hence [17] gets a much more limited length generalization.
In this work, we proposed two inductive scratchpads for the addition task. The inductive scratchpad that requires random spaces in the question can generalize to numbers with 18/20 digits when trained on numbers with up to 10 digits. The other inductive scratchpad based on shifting the operands at each step can generalize to numbers with up to 26 digits when trained on numbers with up to 4 digits at the cost of having a less natural question formatting. Different works use different input formatting, techniques, model sizes, train/test distribution, and evaluation procedures, nonetheless, a rough comparison between different works for the addition task is provided in Table 3 (reported results correspond to the median performance given by different seeds). We believe our solutions for both the parity and addition tasks require one of the least stringent modifications of the input and provide a significant improvement of the length generalization performance compared to prior works even though the models that we have used are often remarkably smaller.
Table 3: Length generalization performance of different methods for the addition task
Work Performance Method and Assumptions
Kazemnejad et al. [34] From 8 digits to 9 digits Using NoPE (no positional embedding)
Shen et al. [17] From 10 digits to 12 digits Scratchpad with recursive format
RASP-L work [35] From 40 digits to 50 digits Reverse order of the output + ‘index hints’ (special tokens before each digit), e.g., $a5b4+a3b7=b1a9$
Jelassi et al. [36] From 5 digits to 15 digits Encoder-only model + relative pos. emb. + padded inputs
Zhou et al. [37] From 40 digits to 65 digits FIRE relative pos. emb. + randomized position encodings + reversed output + index hints
Our random space method From 10 to 18 digits Inductive scratchpad + random space in the input (e.g., $94\_+\_3\_\_1=$)
Our shift method From 4 to 26 digits Inductive scratchpad + random text before each operand (e.g., $fs\$46+ih\$98$)
A.3 Reasoning over graphs
@@
On the modeling front, a group of efforts have primarily focused on baking the appropriate inductive biases into the models for reasoning and logical inference over graphs. The work [81] proposes architectural innovations to use both the context and external knowledge sources. [82] proposes an end-to-end neural model for learning compositional logical rules. The Neurosymbolic approach, presented in [83] involves performing semantic parsing on natural language input using neural components to transform the input into discrete structures suitable for symbolic reasoning by the logic theorem provers. Graphomer [84] extends standard Transformer architecture with a graph reasoning inductive bias. The work of [57] combines Transformers with the inductive bias of the recurrent models. Other works take the route of generating more data using external sources like a Knowledge Graph [85] or extending pre-training [86]. The work in [79] evaluates LLMs’ graph reasoning capabilities in the presence of various prompting techniques like scratchpads/CoT and their variants. This study introduces Build-a-Graph Prompting, an instruction-based prompting method, as well as Algorithmic Prompting, which includes references to relevant graph algorithms in the prompts, as two approaches for enhancing LLMs in solving graph tasks in natural language.
Concurrently with our work, [87] studies the graph reasoning abilities of Transformers. They focus on different tasks such as node count, edge count, connectivity, and shortest path. On the theoretical side, they focus on the expressivity of Transformers. In particular, they show that log-depth Transformers can express the graph connectivity task. This is while in our paper, we focus on learning, showing that for hard enough distributions (i.e., distributions with high globality), regular Transformers cannot learn the connectivity task despite being able to express it. On the empirical side, they show that Transformers can perform well on the connectivity tasks in GraphQA dataset [80], outperforming graph neural networks (GNNs). We note that this is not in contrast to our results. As we saw in Section 1.2 with random graphs, if the input distribution has low globality the connectivity task can become learnable for Transformers.
A.4 Learning measures and lower-bounds for GD
@@ B.1 Implications on random graphs
Figure 7: Performance of a model trained on a balanced distribution of random graphs with 24 nodes and edges where with probability $0.5$ the query nodes are not connected and with probability $0.5$ they are connected and their distance is uniformly selected from $1,2,3,4$. The validation set has the same distribution as the training set showing that the model reaches around $80\%$ accuracy on in-distribution samples. Particularly, the model has perfect accuracy on connected nodes (distance 1-4) and around $60\%$ accuracy on the nodes that are not connected. However, when we tested the model on OOD samples (where some spurious correlations are not present) the model showed a chance level performance. Note that these samples would be of low complexity if the model was actually checking whether there exists a path or not.
@@
We tried to further understand the behavior of this model. By sampling, we computed that one can get an accuracy of around $82\%$ on in-distribution samples just by outputting not-connected if the out-degree of the source query node or the in-degree of the destination query node is zero and connected otherwise. Further, we noticed that this predictor has a high correlation with the output of the model. In particular, in almost all of the cases that the model predicts ‘not-connected’, the source’s out-degree or the destination’s in-degree is zero. (The model may still misclassify some of such samples depending on the random seed.) The latter shows that the model is indeed relying on the degrees of the query nodes as a shortcut.